# Latent Action Diffusion for Cross-Embodiment Manipulation

### v1

@@ Abstract
End-to-end learning approaches offer great potential for robotic manipulation, but their impact is constrained by data scarcity and heterogeneity across different embodiments. In particular, diverse action spaces across different end-effectors create barriers for cross-embodiment learning and skill transfer. We address this challenge through diffusion policies learned in a latent action space that unifies diverse end-effector actions. We first show that we can learn a semantically aligned latent action space for anthropomorphic robotic hands, a human hand, and a parallel jaw gripper using encoders trained with a contrastive loss. Second, we show that by using our proposed latent action space for co-training on manipulation data from different end-effectors, we can utilize a single policy for multi-robot control and obtain up to 13% improved manipulation success rates, indicating successful skill transfer despite a significant embodiment gap. Our approach using latent cross-embodiment policies presents a new method to unify different action spaces across embodiments, enabling efficient multi-robot control and data sharing across robot setups. This unified representation significantly reduces the need for extensive data collection for each new robot morphology, accelerates generalization across embodiments, and ultimately facilitates more scalable and efficient robotic learning.
@@ 1 Introduction
1.1 Contributions
Our contributions are threefold:
  1. 1.
    We introduce a novel framework to learn a cross-end-effector latent action space, which both preserves modality-specific features and cross-modal alignment.
  2. 2.
    We demonstrate the use of our latent action space with latent action diffusion, training capable latent diffusion policies for cross-embodiment manipulation.
  3. 3.
    In real-world experiments, we show that policies trained with cross-embodiment data using our proposed latent action space outperform single-embodiment policies, which indicates the potential of our method to transfer skills across embodiments.
@@ 3.1.2 Retargeting
Figure 3: Qualitative evaluation of the joint latent action space. We encode normalized gripper widths $W\in[0,1]$ (from closed to open) and perform cross-modal reconstruction by decoding them into human hand poses (colored lines on left) and poses for the Faive hand (grey model on right). Existing approaches using retargeting only allow for single-directional retargeting (i.e. human hands to robot hands), which is a limitation our latent action space overcomes. Any modality can be encoded and decoded to any other modality under the alignment constraints of the paired training data.
@@ 3.2 Contrastive Latent Space Learning
3.3 Policy Learning
With a learned latent action space, we can employ any common supervised imitation learning framework such as Diffusion Policy [21] to map shared observations across datasets to latent actions of different embodiments (Fig. 4).
3.3.1 Policy Architecture
We validate our method across two diffusion policy implementations: a transformer-based implementation [22] and a U-Net-based implementation [21]. The architecture for the transformer-based policy is visualized in detail in Figure 4. The transformer-based policy is utilized in experiments with the Faive hand and Franka gripper, whereas the U-Net-based implementation is used for experiments with the mimic hand and the Franka gripper. Both architectures are described in more detail in the appendix in Section 7.1.
3.3.2 Cross-Embodiment Training
For co-training on differently sized datasets, we assign normalized weights $w_{j}$ to all datasets. During training, we seek to combine samples from all datasets to fill batches with $B$ samples in total. We sample per-dataset sub-batches with appropriately rounded sizes $\text{round}(\frac{B}{w_{j}})$, project the actions into the shared latent action space, normalize the sub-batches, and then concatenate them into a single batch for efficient training. Through this mechanism, the weight of each dataset approximately represents a sampling probability for each training step.
4 Experimental Results and Discussion
To validate our proposed framework, we conducted experiments covering three different end-effectors and three tasks across two setups: one with the Faive hand and the Franka gripper and one setup with the mimic hand and the Franka gripper. For each end-effector in each setup, we compare three policies: two single-embodiment policies with and without our proposed latent action space trained on the same data and one cross-embodiment policy co-trained on a data mix that includes additional data from the other end-effector. Additionally, we validate our design choices for our contrastive action model through an ablation study.
4.1 Ablation Study: Contrastive Action Model
@@
4.2 Setup #1: Faive Hand and Franka Gripper
@@ 4.2.1 Performance
4.3 Setup #2: mimic Hand and Franka Gripper
We train policies with data collected using the mimic hand and the Franka gripper for placing a plastic cube inside a box. For each embodiment, we collect 250 episodes of data. The latent space was learned only with data from the mimic hand and the Franka gripper. In addition to an external camera view, we utilize the pose of the Franka arm as observations as well as a wrist camera view for the mimic hand, which is replaced by zero-padding for the Franka gripper. For evaluation, we collect 12 grasps for each embodiment, consisting of 3 tries across 6 grasp locations.
@@ 4.3.1 Performance
Figure 5: Cross-embodiment policy rollouts for two pick and place tasks in different settings. The robots in each setting (left: Franka gripper, Faive hand, right: Franka gripper, mimic hand) are controlled by a single policy, demonstrating multi-robot control.
@@ 4.4 Discussion
Our observations underline the potential of our proposed latent action space for cross-embodiment learning. First, when comparing single-embodiment latent policies to their counterparts with explicit actions, there is no significant performance decrease, which suggests that the learned latent spaces can accurately encode the action spaces of the three embodiments with which we conduct the experiments. Second, the improved average success rate when co-training on data from different end-effectors is a promising sign that the shared, semantically aligned action representations are resulting in helpful inductive biases for the latent policy. These biases lead to improved success rates and reduced standard deviations across different object arrangements, indicating higher robustness and improved spatial generalization. The decrease in performance for the Franka gripper in the mimic-Franka experiment could be attributed to the missing wrist camera during co-training (Section 4.3), which would be in line with the findings from Octo [3] for co-training with missing camera views. An additional factor may also be incidental factors during data collection that could contribute to overall reduced policy performance for the Franka gripper in this setup.
Figure 6: We compare a policy co-trained on cross-embodiment data with single-embodiment policies trained using both latent and explicit actions. For all policies, we show the mean and standard deviation of the number of successful executions across all object locations.
@@ 5 Conclusion
Among current challenges in robotics, enabling effective skill-transfer across diverse embodiments is of crucial importance to both maximize the volume and diversity of suitable training data and to ensure the reusability of training data throughout the lifecycle of different end-effectors. To this end, we frame cross-embodiment learning with different end-effectors as a multimodal representation learning problem and propose a four-stage pipeline to learn capable policies that can control multiple end-effectors. In real-world experiments with the dexterous hands and a Franka parallel gripper, we demonstrate that through co-training on cross-embodiment data with our method for latent action spaces, both multi-robot control and positive skill transfer across embodiments are possible. In particular, the performance improvement of up to 13% indicates that our method facilitates skill transfer between end-effectors with a large embodiment gap and underlines its potential for wider use across a broader range of robot morphologies. Future work includes expanding our method to a more diverse ecosystem of end-effectors and further investigating the behavior of skill transfer across different dataset sizes with more distinct visual differences.

### v2

@@ Abstract
End-to-end learning approaches offer great potential for robotic manipulation, but their impact is constrained by data scarcity and heterogeneity across different embodiments. In particular, diverse action spaces across different end-effectors create barriers for cross-embodiment learning and skill transfer. We address this challenge through diffusion policies learned in a latent action space that unifies diverse end-effector actions. We first show that we can learn a semantically aligned latent action space for anthropomorphic robotic hands, a human hand, and a parallel jaw gripper using encoders trained with a contrastive loss. Second, we show that by using our proposed latent action space for co-training on manipulation data from different end-effectors, we can utilize a single policy for multi-robot control and obtain up to 25% improved manipulation success rates, indicating successful skill transfer despite a significant embodiment gap. Our approach using latent cross-embodiment policies presents a new method to unify different action spaces across embodiments, enabling efficient multi-robot control and data sharing across robot setups. This unified representation significantly reduces the need for extensive data collection for each new robot morphology, accelerates generalization across embodiments, and ultimately facilitates more scalable and efficient robotic learning.
@@ 3.1.2 Retargeting
Figure 3: Qualitative evaluation of the joint latent action space. We encode normalized gripper widths $W\in[0,1]$ (from closed to open) and perform cross-modal reconstruction by decoding them into human hand poses (colored lines on left) and poses for the Faive hand (grey model on right). Existing approaches using retargeting only allow for single-directional retargeting (i.e. human hands to robot hands), which is a limitation our latent action space overcomes. Any modality can be encoded and decoded to any other modality under the alignment constraints of the data.
@@ 3.2 Contrastive Latent Space Learning
3.3 Policy Learning
With a learned latent action space, we employ Diffusion Policy [21] to map shared observations across datasets to latent actions of different embodiments (Fig. 4). We validate our method across two diffusion policy implementations: a transformer-based implementation [22] and a U-Net-based implementation [21]. The architecture for the transformer-based policy is visualized in detail in Figure 4. The transformer-based policy is utilized in experiments with the Faive hand and Franka gripper, whereas the U-Net-based implementation is used for experiments with the mimic hand and the Franka gripper. Both architectures are described in more detail in the appendix in Section 7.2.
@@ 3.3 Policy Learning
4 Experimental Results and Discussion
We conducted experiments covering three different end-effectors and three tasks across two setups: one with the Faive hand and the Franka gripper and one setup with the mimic hand and the Franka gripper. For each end-effector in each setup, we compare single-embodiment policies with one cross-embodiment policy co-trained on data from all end-effectors. Additionally, we validate our design choices for our contrastive action model through an ablation study.
4.1 Ablation Study: Contrastive Action Model
Figure 5: We compare policies co-trained on cross-embodiment data (ours) with single-embodiment policies trained using both latent and explicit actions. For all policies, we show the mean and standard deviation of successful task executions across all object locations. The task setups are shown in Figure 7.
4.2 Setup #1: Pick and Place with Faive Hand and Franka Gripper
@@ 4.2.1 Performance
4.3 Setup #2: Pick and Place with mimic Hand and Franka Gripper
We train policies with data collected using the mimic hand and the Franka gripper for placing a plastic cube inside a box. For each embodiment, we collect 250 episodes of data. The latent space was learned with data from the mimic hand and the Franka gripper. In addition to an external camera view, we utilize the pose of the Franka arm as observations as well as a wrist camera view for the mimic hand, which is replaced by zero-padding for the Franka gripper. For evaluation, we collect 18 grasps for each embodiment, consisting of 3 tries across 6 grasp locations.
@@ 4.3.1 Performance
4.4 Setup #3: Block Stacking with mimic Hand and Franka Gripper
We train policies with data from the mimic hand and the Franka gripper to grasp a block, stack it on top of another block by aligning the pins and then pick up and place the stacked blocks into the box (Figure 6(b)). We collect 200 demontrastions for each embodiment. Both policies use a single external camera and the pose of the Franka arm as observations. For evaluation, we perform 70 rollouts per policy per embodiment.
4.4.1 Performance
We show the completion rates for the individual task stages in Figure 6. For the first stage, a coarse manipulation skill, our method exhibits strong performance with absolute completion rate improvements of 25% and 13%. For the fine-grained manipulation stage, performance slighly decreases for the mimic hand, likely due to the hand partially occluding the object from the camera. For the Franka gripper, the performance improves significantly by 13% and 11%.
@@ 4.5 Discussion
Latent actions accurately encode diverse action spaces. When comparing single-embodiment latent policies to their counterparts with explicit actions, there is no significant performance decrease, which suggests that the learned latent spaces can accurately encode the action spaces of the three embodiments with which we conduct the experiments.
Latent action representations enable multi-robot control and cross-embodiment skill transfer. With our methodology, a single policy can control two highly different end-effectors. Furthermore, latent policies learn helpful shared representations that we observe in improved success rates when co-training on data from different end-effectors for both coarse and fine-grained manipulation tasks. We attribute the performance drop of the Franka gripper in setup #2 to the asymmetric observations: co-training with missing camera views remains an open challenge (similar to [3]). Further, the slight performance decrease for the mimic hand in setup #3 is likely due to the hand occluding the block after grasping it, making fine-grained control more difficult.
In summary, our approach successfully demonstrates that latent action spaces can unify control across diverse robotic embodiments while enabling improved performance through cross-embodiment skill transfer. Learning with asymmetric observation spaces remains remains an important challenge for scaling policy learning.
(a) Task stage completion rates for each task stage for single-embodiment diffusion policy versus cross-embodied latent diffusion policy (ours).
(b) Task stages: i) initial setup ii) picking the first block iii) stacking/inserting it on top of the second block iv) putting both into the box.
Figure 6: We evaluate our methodology across 70 trials per policy per embodiment on a highly challenging block stacking task, comprised of three stages. Both the mimic hand and Franka gripper exhibit a significant performance gain for coarse-grained manipulation (stage 1), whereas performance for fine-grained manipulation is largely improved for the Franka gripper.
@@ 5 Conclusion
Among current challenges in robotics, enabling effective skill-transfer across diverse embodiments is of crucial importance to both maximize the volume and diversity of suitable training data and to ensure the reusability of training data throughout the lifecycle of different end-effectors. To this end, we frame cross-embodiment learning with different end-effectors as a multimodal representation learning problem and propose a four-stage pipeline to learn capable policies that can control multiple end-effectors. In real-world experiments with the dexterous hands and a Franka parallel gripper, we demonstrate that through co-training on cross-embodiment data with our method for latent action spaces, both multi-robot control and positive skill transfer across embodiments are possible. In particular, the performance improvement of up to 25% (average: 13.4%) indicates that our method facilitates skill transfer between end-effectors with a large embodiment gap and underlines its potential for wider use across a broader range of robot morphologies. Future work includes expanding our method to a more diverse ecosystem of end-effectors and further investigating the behavior of skill transfer across different dataset sizes with more distinct visual differences.